OpenStudy (anonymous):

Multiply: 3x+12/8m and m^2/2x+8: A. 3m/16, m =/ 0 B. 3m/16, m =/ 0 and x =/ -4 C. m(x+3)/8, m =/ 0 D. m(x+3)/8, m =/ 0 and x =/-4 Can you help work it out with me?

OpenStudy (anonymous):

sure what do you think you should do first

OpenStudy (anonymous):

\[\frac{3x+12}{8m}*\frac{m^2}{2x+8}\]

OpenStudy (anonymous):

cross multiply?

OpenStudy (anonymous):

i'm sorry i forgot about you haha, how about we factor first ... what can you pull out of the x functions?

OpenStudy (anonymous):

lets say f(x)=3x+12 and g(x)=2x+8 the both have a Common factor

OpenStudy (anonymous):

f(x)=?(x+4) g(x)=?(x+4)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

can you fill in the question marks?